OpenStudy (anonymous):

PLEASE HELP!!!! Marty knows that his living room requires 72 square feet of carpet but is unsure of the dimensions. If he measure the length to be 9 feet, what must its width be? Write only the number rounded to the nearest foot.

Directrix (directrix):

Assumptions: that the living room floor is rectangular and that the carpet is to cover the entire living room floor. A = l*w. 72 = 9*w Divide both sides by 9 to solve for the width.